We need to make 10liters of a 25% solution. We have a 10%, 20%, and a 50% to mix. How many liters of each will be used?
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6
===============
You can use the 50% and the 10%, or the 50% and the 20%
There's no reason to use all 3.